# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ Appendix A Pseudo-stationarity
One can deduce from B.1 that determining the $r$-factorizability of $f$ in (1.4) with $h\in\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$ is equivalent to determining the existence of a feasible solution $(x,g,y,v,\tau)$ with $\|g\|_{\infty}>\lambda$ to (B.12) having nonnegative objective function value. We now turn to B.12 and observe that the objective can be rewritten as a sum with the $i$th summand depending only on $(x_{i},g_{i},y_{\tau(i)},v_{\tau(i)})$. [^6] Specifically, notice that the fourth sum in the objective can be rewritten as $\sum_{i=1}^{m}(Ly_{\tau(i)}+v_{\tau(i)})(\mu y_{\tau(i)}+v_{\tau(i)})$. In addition, from the structure of the constraints in (B.12), we see that the terms $\{(x_{i},g_{i},y_{\tau(i)},v_{\tau(i)})\}_{i\in[m]}$ can be divided into 4 groups depending on whether $i\in[r]$ and $\tau(i)\in[r^{*}]$. These motivate the definitions of the next four associated index sets, for any fixed $\tau\in\mathfrak{P}_{m}$:
$\mathcal{J}_{1}^{\tau}:=[r]\cap\tau^{-1}[r^{*}],~{}\mathcal{J}_{2}^{\tau}:=[r]\setminus\mathcal{J}_{1}^{\tau},~{}\mathcal{J}_{3}^{\tau}:=([m]\setminus[r])\cap\tau^{-1}[r^{*}],~{}\mathcal{J}_{4}^{\tau}:=([m]\setminus[r])\setminus\mathcal{J}_{3}^{\tau}.$ (B.13)
Let $d_{\tau}:=|\mathcal{J}_{2}^{\tau}|$. Then, by the definition of $\{\mathcal{J}_{i}^{\tau}\}_{i\in[4]}$ in B.13, we have
$|\mathcal{J}_{1}|=r-|\mathcal{J}_{2}^{\tau}|=r-d_{\tau},~{}|\mathcal{J}_{3}^{\tau}|=r^{*}-|\mathcal{J}_{1}^{\tau}|=r^{*}-r+d_{\tau},~{}|\mathcal{J}_{4}^{\tau}|=m-r-|\mathcal{J}_{3}^{\tau}|=m-r^{*}-d_{\tau}.$ (B.14)
@@
To solve (B.12), our strategy is to introduce an auxiliary variable $w\in\mathbb{R}$ to transform B.12 to the next equivalent form:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m},w\in\mathbb{R}\end{subarray}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.15)
@@
Problem (B.12) is equivalent to (B.15) in the following sense: for any feasible solution $(x,g,y,v,\tau)$ of (B.12), $(x,g,y,v,\tau,\min_{i\in[r]}x_{i})$ is a feasible solution of (B.12) having the same objective function value; for any feasible solution $(x,g,y,v,\tau,w)$ of (B.15), $(x,g,y,v,\tau)$ is a feasible solution of (B.12) having the same objective function value. Consequently, we have the following result.
Proposition B.2.
There exists a feasible solution $(x,g,y,v,\tau)$ with $\|g\|_{\infty}>\lambda$ to (B.12) having nonnegative objective function value if and only if there exists a feasible solution $(x,g,y,v,\tau,w)$ with $\|g\|_{\infty}>\lambda$ to (B.15) having nonnegative objective function value.
Next, we plan to fix $\tau$ and $w$ to analyze the optimal value and the optimal solution of (B.15). The reason to do so is that the optimization problem can be made separable when $\tau$ and $w$ are fixed. To simplify the calculation, we only consider the case where $\lambda=1$ in the next lemma.
Lemma B.3.
Let $r^{*}\in[m]\cup\{0\}$, $r\in[m]$, and $\infty>L\geq\mu>0$. Let $\tau\in\mathfrak{P}_{m}$ and $w>0$. Let $\{\mathcal{J}_{i}^{\tau}\}_{i\in[4]}$ be defined in B.13. Consider the following optimization problem:
$\sup_{x,g,y,v\in\mathbb{R}^{m}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.16)
@@
Then, the optimization problem (B.16) has optimal solutions, and the optimal value is
$\left(-L\mu|\mathcal{J}_{2}^{\tau}|+|\mathcal{J}_{3}^{\tau}|\frac{L(L-\mu)^{2}}{4\mu}\right)w^{2}.$ (B.17)
@@
  • For all the optimal solutions $(\bar{x},\bar{g},\bar{y},\bar{v})$ of (B.16), we have $\|\bar{g}\|_{\infty}>1$.
  • There exists one optimal solution $(\bar{x},\bar{g},\bar{y},\bar{v})$ of (B.16) such that $\|\bar{g}\|_{\infty}>1$.
@@
$\sum_{i=1}^{m}(Ly_{i}+v_{i})(\mu y_{i}+v_{i})=\sum_{i=1}^{m}(Ly_{\tau(i)}+v_{\tau(i)})(\mu y_{\tau(i)}+v_{\tau(i)}).$ (B.18)
Substituting B.18 into B.16, we get that
$\sup_{x,g,y,v\in\mathbb{R}^{m}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.19)
@@
Observe that (B.19) can be decomposed into the next $m$ subproblems for each $i\in[m]$:
$\sup_{x_{i},g_{i},y_{i},v_{i}\in\mathbb{R}}{}$ $(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.20)
@@
$s.t.{}{}{}{}{}{}$$~{}\begin{cases}x_{i}\geq w,~{}g_{i}=1,~{}y_{\tau(i)}>0,~{}v_{\tau(i)}=1,&~{}~{}\text{if }i\in\mathcal{J}^{\tau}_{1},\\ x_{i}\geq w,~{}g_{i}=1,~{}y_{\tau(i)}=0,~{}v_{\tau(i)}\in[0,1],&~{}~{}\text{if}i\in\mathcal{J}^{\tau}_{2},\\ x_{i}=0,~{}g_{i}\in[0,1+Lw],~{}y_{\tau(i)}>0,~{}v_{\tau(i)}=1,&~{}~{}\text{if }i\in\mathcal{J}^{\tau}_{3},\\ x_{i}=0,~{}g_{i}\in[0,1+Lw],~{}y_{\tau(i)}=0,~{}v_{\tau(i)}\in[0,1],&~{}~{}\text{if }i\in\mathcal{J}^{\tau}_{4},\end{cases}$
where we recall the definition of $\{\mathcal{J}^{\tau}_{i}\}_{i\in[4]}$ in B.13. We now consider the solution and optimal value of each subproblem B.20 for fixed $i$:
@@
    $i\in\mathcal{J}^{\tau}_{1}$. Then B.20 takes the following form:
    $\sup_{x_{i},y_{\tau(i)}}$ $(Lx_{i}+1)(\mu y_{\tau(i)}+1)+(\mu x_{i}+1)(Ly_{\tau(i)}+1)$ (B.21)
@@
    where we used the fact that $g_{i}$ and $v_{\tau(i)}$ are $1$. Denote the objective function of B.21 by $S_{1}$. By direct calculation, we can rewrite $S_{1}(x_{i},y_{\tau(i)})$ as:
@@
    Clearly, the optimal value of (B.21) is 00, and it is achieved if and only if
    $x_{i}=y_{\tau(i)}\geq w.$ (B.22)
@@
    $i\in\mathcal{J}^{\tau}_{2}$. Then B.20 takes the following form:
    $\sup_{x_{i},v_{\tau(i)}}$ $(Lx_{i}+1)v_{\tau(i)}+(\mu x_{i}+1)v_{\tau(i)}-(Lx_{i}+1)(\mu x_{i}+1)-v_{\tau(i)}^{2}$ (B.23)
@@
    where we used the fact that $g_{i}$ and $y_{\tau(i)}$ are $1$ and 00, respectively. Denote the objective function of B.23 by $S_{2}$. By direct calculation, we can rewrite $S_{2}(x_{i},v_{\tau(i)})$ as:
@@
    First, we notice that $S_{2}$ is strictly decreasing on $[0,\infty)$ as a function of $x_{i}$ when $v_{\tau(i)}$ is fixed to be any value in $[0,1]$. This means that
    $\sup_{x_{i}\geq w}S_{2}(x_{i},v_{\tau(i)})=-L\mu w^{2}+(L+\mu)w(v_{\tau(i)}-1)-(v_{\tau(i)}-1)^{2},$ (B.24)
    where the optimal value is achieved if and only if $x_{i}=w$. Let $\tilde{S}$ denote the function on the right hand side of B.24. Then we see $\tilde{S}$ is strictly increasing as a function of $v_{\tau(i)}$ on $(-\infty,1]$ by using the elementary properties of quadratic functions. Therefore, the optimal value of (B.23) is $-L\mu w^{2}$, and it is achieved if and only if
    $x_{i}=w,~{}v_{\tau(i)}=1.$ (B.25)
@@
    $i\in\mathcal{J}^{\tau}_{3}$. Then B.20 has the following form:
    $\sup_{g_{i},y_{\tau(i)}}$ $g_{i}(\mu y_{\tau(i)}+1)+g_{i}(Ly_{\tau(i)}+1)-g_{i}^{2}-(Ly_{\tau(i)}+1)(\mu y_{\tau(i)}+1)$ (B.26)
@@
    where we used the fact that $x_{i}$ and $v_{\tau(i)}$ are 00 and $1$, respectively. Denote the objective function of B.26 by $S_{3}$. By direct calculation we can rewrite $S_{3}$ as follows
@@
    $~{}~{}~{}-L\mu\left(y_{\tau(i)}-\frac{(L+\mu)(g_{i}-1)}{2L\mu}\right)^{2}.$ (B.27)
@@
    Consequently, the optimal value of B.26 is $\frac{L(L-\mu)^{2}w^{2}}{4\mu}$, and is achieved as described in (B.28) and (B.29).
@@
    $i\in\mathcal{J}^{\tau}_{4}$. Then B.20 has the following form:
    $\sup_{g_{i},v_{\tau(i)}}$ $2g_{i}v_{\tau(i)}-g_{i}^{2}-v_{\tau(i)}^{2}$ (B.30)

### v2

@@ 4 Characterization of $r$-factorizability for strongly convex functions
Then, the existence of $h\in\mathfrak{S}(L,\mu,r^{*})$ such that the corresponding $f$ in 1.1 is not $r$-factorizable, is equivalent to the existence of a feasible solution $(x,g,y,v,\tau)$ to 4.16 with $\|g\|_{\infty}>\lambda$ and a nonnegative objective function value.
Proof.
If there exists $h\in\mathfrak{S}(L,\mu,r^{*})$ such that the corresponding $f$ in 1.1 is not $r$-factorizable, then by Proposition 4.1 we know there exists a feasible solution $(x,g,y,v,\tau)$ with $\|g\|_{\infty}>\lambda$ to 4.16 having nonnegative function value. Conversely, if there exists a feasible solution $(x,g,y,v,\tau)$ with $\|g\|_{\infty}>\lambda$ to 4.16 having nonnegative function value, by Lemma 4.2 we may assume that $x$ and $y$ are sorted in descending order, and then by Lemma 4.3 we know there exists $h\in\mathfrak{S}(L,\mu,r^{*})$ such that $f$ in 1.1 is not $r$-factorizable. ∎
Consequently, to prove 1.3, it suffices to solve the optimization problem 4.16 analytically. This is very technically and contains less insight, and hence we defer it to Appendix A. Finally, let us note that 1.3 follows from Proposition 4.4, Proposition A.1 and Proposition A.3.
5 Generalization to the RIP case
When the condition in 1.3 fails, we can find a function $h\in\mathfrak{S}(L,\mu,r^{*})$ such that the corresponding $f$ in 1.1 is not $r$-factorizable. Since $\mathfrak{S}(L,\mu,r^{*})\subseteq\mathfrak{S}(L,\mu,r,r^{*})$, this means the condition in 1.3 is also necessary for the function class $\mathfrak{S}(L,\mu,r,r^{*})$, and it suffices to prove sufficiency in Corollary 1.4. Our proof is largely motivated by the following observation:
Fact 5.1.
If $h\in C^{2}(\mathbb{R}^{m\times n})$ and satisfies 1.7, and the linear subspace $\mathcal{S}\subseteq\mathbb{R}^{m\times n}$ contains only matrices of rank at most $q+r^{*}$, then the restriction $h|_{\mathcal{S}}$ of $h$ on $\mathcal{S}$ satisfies that $h|_{\mathcal{S}}\in C^{2}(\mathcal{S})$ is $\mu$-strongly convex, and $\nabla h|_{\mathcal{S}}$ is $L$-Lipschitz continuous.
Roughly speaking, to prove the sufficiency of the conditions in 1.3 for the function class $\mathfrak{S}(L,\mu,r,r^{*})$, we argue by contradiction. Suppose there exists $h\in\mathfrak{S}(L,\mu,r,r^{*})$ such that the corresponding $f$ in 1.1 is not $r$-factorizable, then there exists a second-order stationary point $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ such that $\bar{X}=\bar{U}\bar{V}^{\top}$ is not a global minimizer of $f$. If we can find a linear subspace $\mathcal{S}$ such that $\mathcal{S}$ contains $\bar{X}$ and $X^{*}$, and $\mathcal{S}$ consists of matrices of rank no more than $r+r^{*}$, then we can apply 1.3 to arrive at a contradiction by restricting $h$ and $f$ on $\mathcal{S}$ and utilizing Fact 5.1. This is achieved by the following lemma.
Lemma 5.2.
Let $A,B\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(A)+\mathrm{rank}(B)=k$. Then, there exists $U\in\mathcal{O}_{m}$ and $V\in\mathcal{O}_{n}$ such that
$A=U\widetilde{\rm Diag}(\sigma(A))V^{\top},\quad B=U\begin{bmatrix}B_{1}&0\\ 0&0\end{bmatrix}V^{\top},$
where $B_{1}\in\mathbb{R}^{k\times k}$.
Proof.
Suppose $\begin{bmatrix}A&B\end{bmatrix}=U_{1}D_{1}V_{1}^{\top}$ and $\begin{bmatrix}A\\ B\end{bmatrix}=U_{2}D_{2}V_{2}^{\top}$ are the corresponding SVDs of $\begin{bmatrix}A&B\end{bmatrix}$ and $\begin{bmatrix}A\\ B\end{bmatrix}$, respectively. Notice that both $\mathrm{rank}\begin{bmatrix}A&B\end{bmatrix}$ and $\mathrm{rank}\begin{bmatrix}A\\ B\end{bmatrix}$ are no more than $k$. Then, the last mkm-k rows of $U_{1}^{\top}A$ and $U_{1}^{\top}B$ are zero because $U_{1}^{\top}\begin{bmatrix}A&B\end{bmatrix}=D_{1}V_{1}^{\top}$, and the last nkn-k columns of $AV_{2}$ and $BV_{2}$ are 0 because $\begin{bmatrix}A\\ B\end{bmatrix}V_{2}=U_{2}D_{2}$. This implies that
$U_{1}^{\top}AV_{2}=\begin{bmatrix}A_{1}&0\\ 0&0\end{bmatrix},\quad U_{1}^{\top}BV_{2}=\begin{bmatrix}B_{1}&0\\ 0&0\end{bmatrix},\quad A_{1},B_{1}\in\mathbb{R}^{k\times k}.$
Let an SVD of $A_{1}$ be given by $A_{1}=U_{3}D_{3}V_{3}^{\top}$. Now, it suffices to take
$U=U_{1}\begin{bmatrix}U_{3}&0\\ 0&I_{m-k}\end{bmatrix}\ \ {\rm and}\ \ V=V_{2}\begin{bmatrix}V_{3}&0\\ 0&I_{n-k}\end{bmatrix}$.
We are now ready to prove in the next proposition the sufficiency of the conditions in 1.3 for the function class $\mathfrak{S}(L,\mu,r,r^{*})$. With Lemma 5.2 in hand, the remaining task is to ensure that we are also able to restrict $F_{r}$ in 1.2 on a corresponding linear subspace to preserve the second-order stationarity of $(\bar{U},\bar{V})$, which is the key ingredient in the proof.
Proposition 5.3.
Assume $h\in\mathfrak{S}(L,\mu,r,r^{*})$ and let $X^{*}\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(X^{*})=r^{*}$ being a global minimizer of $f$ in 1.1. Set $\kappa=\frac{L}{\mu}\geq 1$. Assume that $F_{r}$ in 1.2 has a second-order stationary point $(\bar{U},\bar{V})$ with $\bar{X}:=\bar{U}\bar{V}^{\top}\neq X^{*}$, then both of the following conditions fail:
  • (1)
    $r=m$;
  • (2)
    $r\geq r^{*}$ and $\min\{r,m-r^{*}\}>\frac{(\kappa-1)^{2}}{4}\min\{r^{*},m-r\}.$
Proof.
Let $k=r+r^{*}$. If $k\geq m$, then the result follows immediately from 1.3. Therefore, we assume $k<m$ in the following. Then $r<m$, which proves that $r=m$ is false. Since $k\geq\mathrm{rank}(\bar{X})+\mathrm{rank}(X^{*})$, we can apply Lemma 5.2 to show that there exist $R\in\mathcal{O}_{m}$ and $Q\in\mathcal{O}_{n}$ such that
$R^{\top}X^{*}Q=\begin{bmatrix}X^{*}_{1}&0\\ 0&0\end{bmatrix},\quad R^{\top}\bar{X}Q=\begin{bmatrix}\bar{X}_{1}&0\\ 0&0\end{bmatrix}$,
where $X_{1}^{*},\bar{X}_{1}\in\mathbb{R}^{k\times k}$. Then $\begin{bmatrix}X_{1}^{*}&0\\ 0&0\end{bmatrix}$ is a stationary point of $\tilde{f}$, where $\tilde{f}(W):=f(RWQ^{\top})=\tilde{h}(W)+\lambda\|W\|_{*}$ with $\tilde{h}(W):=h(RWQ^{\top})$. Clearly, we have $\tilde{h}\in\mathfrak{S}(L,\mu,r,r^{*})$ and $\tilde{F}_{r}(U,V):=\tilde{h}(UV^{\top})+\frac{\lambda}{2}\left(\|U\|_{F}^{2}+\|V\|_{F}^{2}\right)=F_{r}(RU,QV)$ for all $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$. In particular, $(R^{\top}\bar{U},Q^{\top}\bar{V})$ is a second-order stationary point of $\tilde{F}_{r}$ with $R^{\top}\bar{U}\bar{V}^{\top}Q=R^{\top}\bar{X}Q\neq\begin{bmatrix}X_{1}^{*}&0\\ 0&0\end{bmatrix}$. Consequently, replacing $f$ by $\tilde{f}$ if necessary, we may assume that
$X^{*}=\begin{bmatrix}X^{*}_{1}&0\\ 0&0\end{bmatrix},\quad\bar{X}=\begin{bmatrix}\bar{X}_{1}&0\\ 0&0\end{bmatrix}$. (5.1)
Let us also note that since $\dim\mathcal{R}(\bar{U})=\mathrm{rank}(\bar{U})=\mathrm{rank}(\bar{V})=\mathrm{rank}(\bar{X})=\dim\mathcal{R}(\bar{X})$ by Remark 3.3(ii), and $\mathcal{R}(\bar{X})\subseteq\mathcal{R}(\bar{U})$, we know that $\mathcal{R}(\bar{U})=\mathcal{R}(\bar{X})$. This implies that $\bar{U}=\begin{bmatrix}\bar{U}_{1}\\ 0\end{bmatrix}$ with $\bar{U}_{1}\in\mathbb{R}^{k\times r}$. Similar arguments on $\bar{X}^{\top}$ show that $\bar{V}=\begin{bmatrix}\bar{V}_{1}\\ 0\end{bmatrix}$ with $\bar{V}_{1}\in\mathbb{R}^{k\times r}$.
Define a function $\widehat{f}:\mathbb{R}^{k\times k}\to\mathbb{R}$ and $\widehat{h}:\mathbb{R}^{k\times k}\to\mathbb{R}$ as:
[…]
$\widehat{f}(K)$ $:=f\left(\begin{bmatrix}K&0\\ 0&0\end{bmatrix}\right)=\widehat{h}(K)+\lambda\left\|\begin{bmatrix}K&0\\ 0&0\end{bmatrix}\right\|_{*}=\widehat{h}(K)+\lambda\|K\|_{*}$
The function $\widehat{f}$ can be viewed as the restriction of $f$ to the linear subspace $\mathcal{S}:=\begin{bmatrix}\mathbb{R}^{k\times k}&0\\ 0&0\end{bmatrix}$. Using Fact 5.1, we know that $\widehat{h}$ is $\mu$-strongly convex, and $\nabla\widehat{h}$ is Lipschitz continuous with modulus $L$. Next, observe that $X_{1}^{*}$ is a global minimizer of $\widehat{f}$ and hence we have $0\in\partial\widehat{f}(X_{1}^{*})$. Since $\widehat{f}(\cdot)=\widehat{h}(\cdot)+\lambda\|\cdot\|_{*}$, we see that $\widehat{f}$ is $\mu$-strongly convex, and hence $X_{1}^{*}$ is the unique stationary point of $\widehat{f}$. Moreover, using 5.1, we know that $\mathrm{rank}(X_{1}^{*})=\mathrm{rank}(X^{*})=r^{*}$. Consequently, $\widehat{h}\in C^{2}(\mathbb{R}^{k\times k})$ satisfies that $\widehat{h}\in\mathfrak{S}(L,\mu,r^{*})$. Next, by direct calculation, the corresponding $\widehat{F}_{r}$ in 1.2 can be written as:
$\widehat{F}_{r}(U,V):=\widehat{h}(UV^{\top})+\frac{\lambda}{2}\left(\|U\|_{F}^{2}+\|V\|_{F}^{2}\right)=h\left(\begin{bmatrix}UV^{\top}&0\\ 0&0\end{bmatrix}\right)+\frac{\lambda}{2}\left(\|U\|_{F}^{2}+\|V\|_{F}^{2}\right),$ (5.2)
which can be viewed as the restriction of $F_{r}$ on the linear subspace $\mathcal{S}_{1}:=\begin{bmatrix}\mathbb{R}^{k\times r}\\ 0\end{bmatrix}\times\begin{bmatrix}\mathbb{R}^{k\times r}\\ 0\end{bmatrix}$. Since $(\bar{U},\bar{V})\in\mathcal{S}_{1}$ is a second-order stationary point of $F_{r}$, we immediately see that $(\bar{U}_{1},\bar{V}_{1})$ is a second-order stationary point of $\widehat{F}_{r}$, since restricting a function on a linear subspace would not change second-order stationarity. The assumption $\bar{X}\neq X^{*}$ gives that $\bar{X}_{1}\neq X^{*}_{1}$ given the representation in 5.1, and hence $\bar{X}_{1}$ is not a global minimizer of $\widehat{f}$, since $\widehat{f}$ is strongly convex and its unique global minimizer is $X^{*}_{1}$. Consequently, $\widehat{f}$ is not $r$-factorizable. By 1.3, we know that both of the following conditions fail:
  • (1’)
    $r=k=r+r^{*}$;
  • (2’)
    $r\geq r^{*}$ and $r>\frac{(\kappa-1)^{2}}{4}r^{*}$.
Since $r+r^{*}=k<m$, we know that $\min\{r,m-r^{*}\}=r$ and $\min\{r^{*},m-r\}=r^{*}$. This proves that (2’) is equivalent to (2), and the proof has been completed. ∎
Appendix A Proof of 1.3
@@ 5 Generalization to the RIP case
In this section, we finish the key step of the proof of 1.3 by solving 4.16. We observe that the objective can be rewritten as a sum with the $i$th summand depending only on $(x_{i},g_{i},y_{\tau(i)},v_{\tau(i)})$. [^10] Specifically, notice that the fourth sum in the objective can be rewritten as $\sum_{i=1}^{m}(Ly_{\tau(i)}+v_{\tau(i)})(\mu y_{\tau(i)}+v_{\tau(i)})$. In addition, from the structure of the constraints in (4.16), we see that the terms $\{(x_{i},g_{i},y_{\tau(i)},v_{\tau(i)})\}_{i\in[m]}$ can be divided into 4 groups depending on whether $i\in[r]$ and $\tau(i)\in[r^{*}]$. These motivate the definitions of the next four associated index sets, for any fixed $\tau\in\mathfrak{P}_{m}$:
$\mathcal{J}_{1}^{\tau}:=[r]\cap\tau^{-1}[r^{*}],~\mathcal{J}_{2}^{\tau}:=[r]\setminus\mathcal{J}_{1}^{\tau},~\mathcal{J}_{3}^{\tau}:=([m]\setminus[r])\cap\tau^{-1}[r^{*}],~\mathcal{J}_{4}^{\tau}:=([m]\setminus[r])\setminus\mathcal{J}_{3}^{\tau}.$ (A.1)
Let $d_{\tau}:=|\mathcal{J}_{2}^{\tau}|$. Then, by the definition of $\{\mathcal{J}_{i}^{\tau}\}_{i\in[4]}$ in A.1, we have
$|\mathcal{J}_{1}|=r-|\mathcal{J}_{2}^{\tau}|=r-d_{\tau},~|\mathcal{J}_{3}^{\tau}|=r^{*}-|\mathcal{J}_{1}^{\tau}|=r^{*}-r+d_{\tau},~|\mathcal{J}_{4}^{\tau}|=m-r-|\mathcal{J}_{3}^{\tau}|=m-r^{*}-d_{\tau}.$ (A.2)
@@
To solve (4.16), our strategy is to introduce an auxiliary variable $w\in\mathbb{R}$ to transform 4.16 to the next equivalent form:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m},w\in\mathbb{R}\end{subarray}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (A.3)
@@
Problem (4.16) is equivalent to (A.3) in the following sense: for any feasible solution $(x,g,y,v,\tau)$ of (4.16), $(x,g,y,v,\tau,\min_{i\in[r]}x_{i})$ is a feasible solution of (4.16) having the same objective function value; for any feasible solution $(x,g,y,v,\tau,w)$ of (A.3), $(x,g,y,v,\tau)$ is a feasible solution of (4.16) having the same objective function value. Consequently, we have the following result.
Proposition A.1.
There exists a feasible solution $(x,g,y,v,\tau)$ with $\|g\|_{\infty}>\lambda$ to (4.16) having nonnegative objective function value if and only if there exists a feasible solution $(x,g,y,v,\tau,w)$ with $\|g\|_{\infty}>\lambda$ to (A.3) having nonnegative objective function value.
Next, we plan to fix $\tau$ and $w$ to analyze the optimal value and the optimal solution of (A.3). The reason to do so is that the optimization problem can be made separable when $\tau$ and $w$ are fixed. To simplify the calculation, we only consider the case where $\lambda=1$ in the next lemma.
Lemma A.2.
Let $r^{*}\in[m]\cup\{0\}$, $r\in[m]$, and $\infty>L\geq\mu>0$. Let $\tau\in\mathfrak{P}_{m}$ and $w>0$. Let $\{\mathcal{J}_{i}^{\tau}\}_{i\in[4]}$ be defined in A.1. Consider the following optimization problem:
$\sup_{x,g,y,v\in\mathbb{R}^{m}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (A.4)
@@
Then, the optimization problem (A.4) has optimal solutions, and the optimal value is
$\left(-L\mu|\mathcal{J}_{2}^{\tau}|+|\mathcal{J}_{3}^{\tau}|\frac{L(L-\mu)^{2}}{4\mu}\right)w^{2}.$ (A.5)
@@
  • For all the optimal solutions $(\bar{x},\bar{g},\bar{y},\bar{v})$ of (A.4), we have $\|\bar{g}\|_{\infty}>1$.
  • There exists one optimal solution $(\bar{x},\bar{g},\bar{y},\bar{v})$ of (A.4) such that $\|\bar{g}\|_{\infty}>1$.
@@
$\sum_{i=1}^{m}(Ly_{i}+v_{i})(\mu y_{i}+v_{i})=\sum_{i=1}^{m}(Ly_{\tau(i)}+v_{\tau(i)})(\mu y_{\tau(i)}+v_{\tau(i)}).$ (A.6)
Substituting A.6 into A.4, we get that
$\sup_{x,g,y,v\in\mathbb{R}^{m}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (A.7)
@@
Observe that (A.7) can be decomposed into the next $m$ subproblems for each $i\in[m]$:
$\sup_{x_{i},g_{i},y_{i},v_{i}\in\mathbb{R}}$ $(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (A.8)
@@
[…]
where we recall the definition of $\{\mathcal{J}^{\tau}_{i}\}_{i\in[4]}$ in A.1. We now consider the solution and optimal value of each subproblem A.8 for fixed $i$:
@@
    $i\in\mathcal{J}^{\tau}_{1}$. Then A.8 takes the following form:
    $\sup_{x_{i},y_{\tau(i)}}$ $(Lx_{i}+1)(\mu y_{\tau(i)}+1)+(\mu x_{i}+1)(Ly_{\tau(i)}+1)$ (A.9)
@@
    where we used the fact that $g_{i}$ and $v_{\tau(i)}$ are $1$. Denote the objective function of A.9 by $S_{1}$. By direct calculation, we can rewrite $S_{1}(x_{i},y_{\tau(i)})$ as:
@@
    Clearly, the optimal value of (A.9) is 0, and it is achieved if and only if
    $x_{i}=y_{\tau(i)}\geq w.$ (A.10)
@@
    $i\in\mathcal{J}^{\tau}_{2}$. Then A.8 takes the following form:
    $\sup_{x_{i},v_{\tau(i)}}$ $(Lx_{i}+1)v_{\tau(i)}+(\mu x_{i}+1)v_{\tau(i)}-(Lx_{i}+1)(\mu x_{i}+1)-v_{\tau(i)}^{2}$ (A.11)
@@
    where we used the fact that $g_{i}$ and $y_{\tau(i)}$ are $1$ and 0, respectively. Denote the objective function of A.11 by $S_{2}$. By direct calculation, we can rewrite $S_{2}(x_{i},v_{\tau(i)})$ as:
@@
    First, we notice that $S_{2}$ is strictly decreasing on $0,\infty)$ as a function of $x_{i}$ when $v_{\tau(i)}$ is fixed to be any value in $[0,1]$. This means that
    $\sup_{x_{i}\geq w}S_{2}(x_{i},v_{\tau(i)})=-L\mu w^{2}+(L+\mu)w(v_{\tau(i)}-1)-(v_{\tau(i)}-1)^{2},$ (A.12)
    where the optimal value is achieved if and only if $x_{i}=w$. Let $\tilde{S}$ denote the function on the right hand side of A.12. Then we see $\tilde{S}$ is strictly increasing as a function of $v_{\tau(i)}$ on $(-\infty,1]$ by using the elementary properties of quadratic functions. Therefore, the optimal value of (A.11) is $L\mu w^{2}$, and it is achieved if and only if
    $x_{i}=w,~v_{\tau(i)}=1.$ (A.13)
@@
    $i\in\mathcal{J}^{\tau}_{3}$. Then A.8 has the following form:
    $\sup_{g_{i},y_{\tau(i)}}$ $g_{i}(\mu y_{\tau(i)}+1)+g_{i}(Ly_{\tau(i)}+1)-g_{i}^{2}-(Ly_{\tau(i)}+1)(\mu y_{\tau(i)}+1)$ (A.14)
@@
    where we used the fact that $x_{i}$ and $v_{\tau(i)}$ are 0 and $1$, respectively. Denote the objective function of A.14 by $S_{3}$. By direct calculation we can rewrite $S_{3}$ as follows
@@
    $~~~-L\mu\left(y_{\tau(i)}-\frac{(L+\mu)(g_{i}-1)}{2L\mu}\right)^{2}.$ (A.15)
@@
    Consequently, the optimal value of A.14 is $\frac{L(L-\mu)^{2}w^{2}}{4\mu}$, and is achieved as described in (A.16) and (A.17).
@@
    $i\in\mathcal{J}^{\tau}_{4}$. Then A.8 has the following form:
    $\sup_{g_{i},v_{\tau(i)}}$ $2g_{i}v_{\tau(i)}-g_{i}^{2}-v_{\tau(i)}^{2}$ (A.18)